3x-2y=12
6x-4y=14

Answer by Alan3354(69443)

3x-2y=12
6x-4y=14 --> 3x-2y=7
--------------------
3x-2y=12
3x-2y=7
-----------
3x-2y can't be both 12 & 7. --> inconsistent.
If you graph them, the lines are parallel.
No intersection, no solution.